20 g of an ideal gas contains only atoms of S and O occupies 5.6 L at 1 atm and 273 K. What is the mol. wt. of gas?
1
Expert's answer
2020-11-19T08:51:16-0500

5.6 Liter of gas at 1atm and 273 k .


Moles of gas = 5.6 / 22.4 = 1/4 = 0.25 moles.


20 g of gas .


Moles of gas = 20 / M .


So, moles of gas = 0.25 = 20 / M .

M = 20 / 0.25 = 80 .


So , gas is SO3 ..